OpenStudy (jensine):

sovle the equation (x+7)(x+3)=0

OpenStudy (anonymous):

x = -7 x = -3

OpenStudy (anonymous):

cuz (-7+7)(x+3) = 0 and (x+7)(-3+3)=0

OpenStudy (anonymous):

To solve it you would set each part equal to 0 and solve. So you'd do x+7=0 and x+3=0 And then solve for x for each. That's how you do it for when you have another instance of that question.